OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s). -x^2 - 4x; x = –3 A. 3 B. –1 C. 11 D. –17

OpenStudy (anonymous):

careful use of parentheses \[-(-3)^2-4\times (-3)=-9+12=3\]

OpenStudy (anonymous):

you have to replace -3 in each statement:\[-(-3)^{2}-4(-3)\].. the result is 3

OpenStudy (anonymous):

thank you both very much i didn't put parentheses , that's why i never got any answer choice